OpenStudy (anonymous):

Determine which of the following point(s) lie on the graph y= square root of x+2 a. (-6,4) b. (7,-3) c. a&b d. neither

OpenStudy (anonymous):

\[y = \sqrt{x + 2}\]Just plug in the coordinate choices with their respective variables

OpenStudy (jusaquikie):

like he said, the numbers are coordinates (x,y) so \[4=\sqrt{-6+2}\] this is square root of -4 which is imaginary \[-3=\sqrt{7+2}\] this is 3 not negative 3 so this one is a no so i would go with D none of the above

OpenStudy (skullpatrol):